a cleaning solution has a volume of 0.4 L. in the cleaning solution is 0.2 moles of detergent. what is the molar concentration of the detergent in the solution?


1
Expert's answer
2021-01-13T13:55:50-0500

M=n/V=0.2/4=0.05


Answer: 0.05M